OpenStudy (anonymous):

What is the approximate area of the shaded region? A. 178.5 square centimeters B. 100 square centimeters C. 78.5 square centimeters D. 21.5 square centimeters

OpenStudy (harsha19111999):

arear = pi*r^2

OpenStudy (harsha19111999):

Can you find it out?

OpenStudy (anonymous):

Area of square - area of circle = area of shaded region...

OpenStudy (dan815):

ahem.,,

OpenStudy (harsha19111999):

\[\pi(10)^2\]

OpenStudy (anonymous):

d= 10 so radius is 5..

OpenStudy (harsha19111999):

Its 68.6

OpenStudy (anonymous):

\[(10^2) - (\pi5^2) = area of shaded region\] solve...

OpenStudy (anonymous):

100 - 25pi...

OpenStudy (anonymous):

C.?

OpenStudy (anonymous):

it's 21.5...

OpenStudy (anonymous):

100 - 25pi = 100 - 25(3.14) = 1000 - 78.5 = 21.5 D.

OpenStudy (harsha19111999):

Option D